OpenStudy (anonymous):

Find the volume of a square pyramid with base edges of 32 cm and a slant height of 30 cm.

OpenStudy (jmark):

Given Base edge and slant height (l) of pyramid as 32 , 30cm. To find the height of Pyramid we use h= √(l2-b2) = √(302-162) (half of the base value is taken) = √644 = 25.5cm Assign the value of base and height value in below formula Volume of Square Pyramid = 1/3 x b2xh = 1/3 x322 x 25.5 = 8704cm3 ICSE board papers - http://icse.edurite.com/icse-sample-papers.html